Check my work Comparative Statements of Shareholders' Equity for Locke Intertechnology Corporation were reported as follows for the fiscal years ending December 31, 2016, 2017, and 2018. Statements of Shareholders' Equity For the Years Ended Dec. 31, 2016, 2017, and 2018 (s in millions) Preferred Stock, Common Stock, $1 Additional Paid-in Retained Sharehol Balance at January 1, 2016 $ 65S $ 575 570 90 1,910 Sale of preferred shares Sale of common shares, 7/1 Cash dividend, preferred Cash dividend, conmon Net income $2,550 577 96 (1 (18 400 3,604 (81 400 (30) (22) Balance at December 31, 2016 1,235 (45) 2,291 Retirement of common shares, 4/1 Cash dividend, preferred Cash dividend, common 3-for-2 split effected in the form of a (22 common stock dividend, 8/12 Net income 32.5 (32.5) 500 500 4,000 Balance at December 31, 2017 97.5 9.75 1,157.5 . 2,738 (99.75) 101 common stock dividend, 5/1 Sale of common shares, 9/1 Cash dividend, preferred Cash dividend, common Net income 90 36 39 (2 (38 (38) Balance at December 31, 2018 $110.25 $1,283.5 3,042-25 $4,443 Required: Infer from the statements the events and transactions that affected Locke Intertechnology Corporation's shareholders' equity and compute earnings per share as it would have appeared on the income statements for the years ended December 31, 2016, 2017, and 2018. No potential common shares were outstanding during any of the periods shown. (Enter your answers in millions rounded to 2 decimal places (ie., 5,500,000 should be entered as 5.50).) K Prev 9 of 9 Next
Help Save & 97.5 75 Balance at December 31,Z01Y L157. ,000 I0s common stock dividend, 5/1 Sale of common shares, 9/1 Cash dividend, prefer Cash dividend, common Net incone 90 36 (99.75) (38) $3,042.25 39 (2 (38 Balance at December 31, 2018 $110.25 $1,283.5 $4,443 Required: Infer from the statements the events and transactions that affected Locke Intertechnology Corporation's shareholders' equity and compute earnings per share as it would have appeared on the income statements for the years ended December 31,2016,2017, and 2018. No potential common shares were outstanding during any of the periods shown. (Enter your answers in millions rounded to 2 decimal places (i.e., 5,500,000 should be entered as 5.50).) 2016 2017 2018 K Prev 9 f Next
Answer #1

numerator = net income - preferred dividend

denominator = number of outstanding common shares

2016

numeratio = 400-1 = 399

denominator = 65-(6*(6/12)) = 62

EPS = 399/62 = 6.44

2017

numeratio = 500-1 = 499

denominator = (71*1.50)-((6*(6/12)*1.50) = 102

EPS = 499/102 = 4.89

2018

numeratio = 444-2 = 442

denominator = (97.5*1.10)+((3*(4/12))= 108.25

EPS = 442/108.25 = 4.08

therefore,

year

Numerator

/

Denominator

=

Earnings per share

2016

399

/

62

=

6.44

2017

499

/

102

=

4.89

2018

442

/

108.25

=

4.08